Three solutions X,Y,Z of HCL are mixed to produce 100 ml of 0.1 M solution. The molarities of X, Y, Z are 0.07 M, 0.12 M and 0.15 M respectively. What respective volumes of X, Y and Z should be mixed?

Grade:

2 Answers

Arun
4 years ago
Dear student
Β 
I am getting my answer as 55ml, 20 ml, 25 ml.
Please let me know of this is correct then I will share my solution.
Khimraj
4 years ago
1. The amount of HCl in the resulting solution is the following: 𝑐𝑐(HCl) = 𝑛𝑛(HCl) 𝑉𝑉(solition) Consequently 𝑛𝑛(HCl) = 𝑐𝑐(HCl) Β· 𝑉𝑉(solition) = 100 π‘šπ‘šπ‘šπ‘š 1000 Β· 0.1 π‘šπ‘šπ‘šπ‘šπ‘šπ‘š 𝐿𝐿 = 0.01 π‘šπ‘šπ‘šπ‘šπ‘šπ‘š 2. The volume of HCl in every solution is marked with x, y and z. We can write the equation: 𝑉𝑉Σ(HCl) = 𝑉𝑉π‘₯π‘₯(HCl) + 𝑉𝑉𝑦𝑦(HCl) + 𝑉𝑉𝑧𝑧(HCl) = π‘₯π‘₯ + 𝑦𝑦 + 𝑧𝑧 = 0.1 𝐿𝐿 Or π‘₯π‘₯ + 𝑦𝑦 + 𝑧𝑧 = 0.1 3. Then we write the formula for total amount of HCl: 𝑛𝑛Σ(HCl) = 𝑛𝑛π‘₯π‘₯(HCl) + 𝑉𝑉𝑦𝑦(HCl) + 𝑛𝑛𝑧𝑧(HCl) = 𝑐𝑐Σ(HCl) Β· 𝑛𝑛Σ(HCl) = 𝑐𝑐π‘₯π‘₯(HCl) Β· 𝑛𝑛π‘₯π‘₯(HCl) + 𝑐𝑐𝑦𝑦(HCl) Β· 𝑛𝑛𝑦𝑦(HCl) + 𝑐𝑐𝑧𝑧(HCl) Β· 𝑛𝑛𝑧𝑧(HCl) = 0.07π‘₯π‘₯ + 0.12𝑦𝑦 + 0.15𝑧𝑧 = 0.01 π‘šπ‘šπ‘šπ‘šπ‘šπ‘š Or 0.07π‘₯π‘₯ + 0.12𝑦𝑦 + 0.15𝑧𝑧 = 0.01 Or 7π‘₯π‘₯ + 12𝑦𝑦 + 15𝑧𝑧 = 1 4. The elimination of variable x it can be expressed by y, that is the solution Z is substituted by Y by way of coefficient expressed by ratio of respective molarities: 𝑐𝑐𝑧𝑧(HCl) 𝑐𝑐𝑦𝑦(HCl) = 1.25 Consequently, 𝑧𝑧 = 1.25𝑦𝑦 5. The set of equations can be composed: οΏ½ π‘₯π‘₯ + 𝑦𝑦 + 𝑧𝑧 = 0.1 7π‘₯π‘₯ + 12 + 15𝑧𝑧 = 1 If 𝑧𝑧 = 1.25𝑦𝑦, then οΏ½ π‘₯π‘₯ + 𝑦𝑦 + 1.25𝑦𝑦 = 0.1 7π‘₯π‘₯ + 12𝑦𝑦 + 18.75𝑦𝑦 = 1 Then οΏ½ π‘₯π‘₯ + 2.25𝑦𝑦 = 0.1 7π‘₯π‘₯ + 30.75𝑦𝑦 = 1 οΏ½ π‘₯π‘₯ = 0.1 βˆ’ 2.25𝑦𝑦 7π‘₯π‘₯ + 30.75𝑦𝑦 = 1 Then 7 Β· (0.1 βˆ’ 2.25𝑦𝑦) + 30.75𝑦𝑦 = 1 0.7 βˆ’ 15.75𝑦𝑦 + 30.75𝑦𝑦 = 1 15𝑦𝑦 = 0.3 𝑦𝑦 = 0.02 𝑧𝑧 = 1.25𝑦𝑦 = 0.025 π‘₯π‘₯ = 0.1 βˆ’ 𝑦𝑦 βˆ’ 𝑧𝑧 = 0.1 βˆ’ 0.02 βˆ’ 0.025 = 0.055 6. The volumes of solutions X, Y, Z are: 𝑉𝑉π‘₯π‘₯(HCl) = 0.055 𝐿𝐿 = 55 π‘šπ‘šπ‘šπ‘š 𝑉𝑉𝑦𝑦(HCl) = 0.025 𝐿𝐿 = 25 π‘šπ‘šπ‘šπ‘š 𝑉𝑉𝑧𝑧(HCl) = 0.020 𝐿𝐿 = 20 π‘šπ‘šπ‘šπ‘š Answer: The volumes of X, Y, Z solutions are 0.055 L, 0.025 L and 0.020 L respectively.